### v1

@@ 4.2 Enhancing Training with Label Descriptions
The generated descriptions are then processed by a pre-trained BERT model to obtain their corresponding description representations. For task $t$, each relation $r\in\mathcal{R}_{t}$ is associated with a set of label description representations $Des_{r}=\{{\bm{d}}_{r,i}\in\mathbb{R}^{d}\ |\ i=1,\dots,D\}$. Given an input ${\bm{x}}$, our objective is to train prompts that guide the input to the correct relation $y$. We aim to pull $f_{\theta}({\bm{x}}_{p})$ closer to the label description representation of the correct relation while pushing it away from the representations of other labels. To achieve this, we introduce a contrastive loss that enforces clearer separation between the target label and irrelevant labels:
@@
where $Des_{y}$ is the set of description representations for the true label $y$. This loss ensures that the representation of the input sample is strongly associated with its corresponding label while reducing its association with incorrect labels, thereby enhancing the model’s discriminative power. Consequently, incorporating label descriptions improves the stability and reliability of learned representations, ensuring better performance even when faced with previously unseen or misclassified samples.
@@
For each new task $\mathcal{T}_{t}$, a corresponding prompt pool $\mathcal{P}_{t}$ is created. During each training step, as described in the aforementioned strategy, $K$ prompts are selected, and the associated prompted embedding feature, denoted as ${\bm{x}}_{p}$, is fed into the pre-trained transformer encoder $f_{\theta}$ and the final classifier $g_{\phi}$. The overall objective can be expressed as follows:
@@
where $\alpha$ and $\beta$ are hyperparameters. The first term in Equation (19) represents the classification loss. Importantly, only the prompt parameters in the current task’s prompt pool $\mathcal{P}_{t}$ and the final classifier $g_{\phi}$ are updated during task $t$ training. The pre-trained BERT model and all previous prompt pools $\mathcal{P}_{1},\dots,\mathcal{P}_{t-1}$ remain frozen. By incorporating task-specific prompt pools and label descriptions, we propose a comprehensive training framework that enhances the model’s ability to accurately predict relations within each task. This approach improves within-task prediction performance, thereby contributing to better continual learning performance as shown in Equation (1). In the following section, we introduce our strategy to further enhance task-identity inference performance.
@@ 4.3 Cascade Voting for Task Prediction
Input: Voting results $\mathcal{V}^{0}({\bm{x}})$ and $\mathcal{V}^{1}({\bm{x}})$ of prompt pool $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, Maximum allowed experts $m$, Confidence scores $\{\text{Score}^{i}_{t}({\bm{x}})\ |\ 2\leq i\leq k-1,i\leq t\leq k\}$
@@
7:  for $i\leftarrow start$ to $end$ do
8:     Determine $\mathcal{V}^{i}({\bm{x}})$ via Equation (23)
@@
During training, the model is provided with a dataset corresponding to a specific task and optimizes the associated task-specific prompt pool. However, during inference, a key challenge arises: the model no longer receives explicit task labels for incoming inputs. Consequently, it must autonomously determine which task-specific prompt pool to employ for generating accurate predictions.
To address this, HiDe-Prompt (Wang et al., 2024) introduces an additional MLP head $\hat{g}_{\psi}$, as a task predictor. This head leverages the pre-trained unprompted feature $f_{\theta}({\bm{x}})$ to predict the task label. However, as discussed in Section 3, treating each task as a distinct class can be suboptimal because tasks are defined by their order of appearance rather than by meaningful semantic differences. In response, Le et al. (2025a) increase the predictor’s output dimension to match the number of encountered relations, then derives the task label from the relation it classifies. Despite this, it still relies on pre-trained knowledge and overlooks the potential task-identification capabilities of the experts with specialized knowledge $\mathcal{P}_{t}$.
In contrast, our approach leverages Cascade Voting (Zhou et al., 2024) to determine task identity. By discarding the MLP-based classifier head, Cascade Voting eliminates the need for explicit training for task prediction. Instead, it employs a voting strategy in which the specialized knowledge of experts in each prompt pool $\mathcal{P}_{t}$ is used collaboratively to determine the task identity for a given instance.
Formally, for an instance ${\bm{x}}$ of relation $r$, let ${\bm{z}}_{r}^{i}=f_{\theta,\mathcal{P}_{i}}({\bm{x}})\in\mathbb{R}^{d}$ be the representation obtained by passing ${\bm{x}}$ through the pre-trained transformer combined with prompt pool $\mathcal{P}_{i}$. We assume during the $t$-th task, the class-conditional distribution of $\mathcal{D}^{i}_{r,t}=\{{\bm{z}}_{r}^{i}=f_{\theta,\mathcal{P}_{i}}({\bm{x}})|({\bm{x}},r)\in\mathcal{D}_{t}\}$ follows a multivariate Gaussian distribution $\mathcal{N}(\bm{\mu}^{i}_{r,t},\bm{\Sigma}^{i}_{t})$ for each prompt pool $\mathcal{P}_{i}$ trained up to task $t$ and for all relations $r\in\mathcal{R}_{t}$, where the covariance $\bm{\Sigma}^{i}_{t}$ is shared by all relations within the task. Based on this assumption, we use the maximum likelihood estimator to approximate the mean and covariance as follows:
@@
Because continual relation extraction precludes access to data from previously encountered tasks, no distribution can be established for how prompt pool $\mathcal{P}_{i}$ affects any earlier task $j$ with $j<i$. To overcome it, we employ cascade voting to determine the appropriate prompts for relation classification during inference. This approach comprises two stages. First, each prompt pool, conceptualized as a specialized set of experts for a given task, independently assesses and votes on the suitability of its prompts. Subsequently, votes from all prompt pools are aggregated to identify the most appropriate prompt pool for the given instance.
Voting Procedure of a Prompt Pool
The voting procedure of a prompt pool $\mathcal{P}_{i}$ begins by passing the input ${\bm{x}}$ through the pre-trained backbone $f_{\theta}$ combined with the prompt pool $\mathcal{P}_{i}$ to obtain the feature ${\bm{z}}^{i}=f_{\theta,\mathcal{P}_{i}}({\bm{x}})$. Given that each relation $r\in\mathcal{R}_{t}$ is characterized by stored multivariate Gaussian distributions $\mathcal{N}(\bm{\mu}^{i}_{r,t},\bm{\Sigma}^{i}_{t})$, we compute the voting score of $\mathcal{P}_{i}$ for each task $t\geqslant i$. In particular, we measure the Mahalanobis distance between ${\bm{z}}^{i}$ and each Gaussian $\{\mathcal{N}(\bm{\mu}^{i}_{r,t},\bm{\Sigma}^{i}_{t})\ |\ r\in\mathcal{R}_{t}\}$, thereby producing $\mathcal{P}_{i}$’s decision for task $t$ as follows:
@@
Finally, the outcome task identity predicted by $\mathcal{P}_{i}$ is determined by:
@@
As the range of tasks covered by each prompt pool can vary, we adopt a cascade voting mechanism to mitigate boundary mismatches across prompt pools. We first employ the pre-trained BERT model $\mathcal{P}_{0}$ as the initial expert. Since both $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ can predict any subsequent task, their voting results are considered first. If $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ concur on a task label, that label is immediately assigned to ${\bm{x}}$.
If $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ disagree, we then gather votes from all relevant prompt pools $\{\mathcal{P}_{i}\ |\ i\in[start,end]\}$, where $start$ and $end$ derive from the ranges indicated by the disagreement between the two initial predictors. The task identity that receives the majority of votes is assigned to ${\bm{x}}$, and the corresponding prompt pool is selected for subsequent relation classification. This cascade voting strategy thus limits the number of prompt pools involved at each decision step and effectively handles inconsistencies in prediction ranges. A detailed algorithmic description is provided in Algorithm 1. Our approach enhances task-identity prediction accuracy, as demonstrated in Section 5.3, leading to improved TII performance and, consequently, more effective continual relation extraction, as described by Equation (1).
Comparison to MLP-based Approach
Both the MLP-based head strategy proposed by Le et al. (2025a) and the cascade voting mechanism introduced by Zhou et al. (2024) share the common principle of treating each observed relation independently, rather than grouping all relations within a single task into a unified category. In the MLP-based paradigm, each head corresponds to a specific relation, while in the cascade voting framework, every relation within a task is scored and a single highest-scoring relation is selected as that task’s representative. Subsequently, the representatives from all tasks are assessed together, and the task yielding the highest score is designated as the final classification decision.
A key distinction lies in whether task classification is explicitly trained. Under the MLP-based paradigm, a standard supervised task-prediction approach is employed: each task is allocated its own MLP head, and the classification result is given by the head returning the highest probability. In contrast, cascade voting dispenses with the MLP classifier, instead leveraging inference-time computations (e.g., distance-based metrics) to identify the most appropriate task.
Regarding memory usage, the MLP-based approach only requires storing relation-specific distributions to facilitate data reconstruction and mitigate catastrophic forgetting. The cascade voting mechanism, however, also relies on relation distributions that are further specialized by each prompt pool, resulting in a larger overall number of stored distributions compared to the MLP-based counterpart. Despite this increased memory demand, a notable advantage of cascade voting is that it eliminates additional training overhead, thereby expediting the overall training process compared to MLP-based classifiers.
Figure 4: Data flow diagram. First, the task identity of the input ${\bm{x}}$ is inferred via cascade voting, which determines the appropriate prompt pool. The input ${\bm{x}}$ then queries this prompt pool to retrieve prompts whose keys most closely match the query $q({\bm{x}})$. The selected prompt is prepended to the embedded input ${\bm{x}}_{e}$, yielding the prompted input ${\bm{x}}_{p}$. This prompted input is fed into the BERT encoder, from which the embeddings at the positions of the entities $E_{1}$ and $E_{2}$ are extracted and concatenated. Finally, this concatenated embedding is passed to the relation classifier $g_{\phi}$, which predicts the relation label $y$ of the original input ${\bm{x}}$.
@@ 4.4 Mitigating Classification Bias with Generative Models
A key element of our framework is the MLP head relation classifier $g_{\phi}$. To mitigate catastrophic forgetting, we retain the representations of previously learned relations by deploying a generator that encodes their input distribution. Specifically, we employ Gaussian distributions $\mathcal{N}(\bm{\mu}_{r}^{t},\bm{\Sigma}^{t})$ for each relation $r\in\mathcal{R}_{t}$, where $\bm{\mu}_{r}^{t}=\bm{\mu}^{t}_{r,t}$ and $\bm{\Sigma}^{t}=\bm{\Sigma}^{t}_{t}$. These are computed from the prompt pool $\mathcal{P}_{t}$ of the current task as described in Section 4.3.
By using a prompt-based generative model, the representations of previously learned relations can be reconstructed and replayed for subsequent tasks, thereby preserving essential relational knowledge. For each relation, storage requirements are minimal: only a mean vector and a covariance matrix are kept, significantly reducing memory overhead while preserving learned features. Furthermore, this design effectively addresses the issue of catastrophic forgetting by eliminating the need to store observed instances explicitly. Instead, the generative models capture and reproduce the distribution of the input space, enabling the model to simulate previously encountered relations without directly accessing historical data. This approach not only enhances computational efficiency but also ensures compliance with privacy constraints, as raw data is never retained.
These generative models produce a representation set ${\bm{z}}_{r}$ for each relation $r$, which is then used to train the relation classifier. The MLP classifier $g_{\phi}$ is trained by minimizing the cross-entropy loss as below:
$\mathcal{L}(\phi)=\sum_{i=1}^{t}\sum_{r\in\mathcal{R}_{i}}\sum_{{\bm{z}}_{r}\sim\mathcal{N}(\bm{\mu}_{r}^{t},\bm{\Sigma}^{t})}-\mathrm{log}\frac{\exp(g_{\phi}({\bm{z}}_{r})[r])}{\sum_{r^{\prime}\in\widehat{\mathcal{R}}_{t}}\exp(g_{\phi}({\bm{z}}_{r})[r^{\prime}])}.$ (24)
For a comprehensive overview of the training procedure, each phase is detailed in Algorithm 2. The data flow diagram illustrating the inference process is provided in Figure 4.
@@
2:  for $epoch\leftarrow 1$ to $training\_epoch$ do
@@
7:  Update ^$\hat{\mathcal{R}}_{t}\leftarrow\hat{\mathcal{R}}_{t-1}\cup\mathcal{R}_{t}$
@@
5 Experiments
5.1 Experimental Settings
Datasets
To ensure consistency and fairness in our evaluation, we follow the protocols outlined by Zhou et al. (2024); Le et al. (2025a) and assess the performance of WAVE++ on two widely used CRE datasets:
  • 1.
    TACRED (Zhang et al., 2017) comprises 106,264 instances spanning 41 relation types. instances spanning 41 relation types. Following the experimental setup of Cui et al. (2021), we partition this dataset into 10 distinct subsets.
  • 2.
    FewRel (Han et al., 2018) contains 56,000 instances encompassing 80 relation types. In accordance with Wang et al. (2019), we divide this dataset into 10 subsets.
Baseline Models
We compare WAVE++ with various prompt-based methods, including L2P (Wang et al., 2022b), HiDe-Prompt (Wang et al., 2024), EoE (Zhou et al., 2024) and WAVE-CRE (Le et al., 2025a). Notably, WAVE-CRE (Le et al., 2025a) is a simplified version of WAVE++. While WAVE-CRE also employs task-specific prompt pools, WAVE++ extends this strategy by integrating label descriptions to better capture relational information, as well as replacing MLP classifier with cascade voting for task identity prediction.
We further contrast our approach with traditional rehearsal-based methods: RP-CRE (Cui et al., 2021), ACA (Wang et al., 2022b), CRL (Zhao et al., 2022), CDec (Xia et al., 2023), CEAR (Zhao et al., 2023), RationaleCL (Xiong et al., 2023), CREST (Le et al., 2024c), and DP-CRE (Huang et al., 2024). For all methods, we utilize a pre-trained BERT encoder (Devlin, 2018) as the backbone architecture. We report model performance in terms of average accuracy across five distinct random runs.

### v2

@@ 4.1 Task-specific Prompt Pool
4.2 Enhancing Training with Label Descriptions
Let ${\bm{x}}_{e}$ be the embeddings of an input sequence $\bm{x}$}. After incorporating prompts via prefix-tuning, the modified input embeddings are denoted as ${\bm{x}}_{p}$. The model’s output is then given by $g_{\phi}(f_{\theta}({\bm{x}}_{p}))$, where $f_{\theta}$ is the pre-trained transformer encoder and $g_{\phi}$ is the final relation classifier, parameterized by $\phi$.
In our framework, each task is associated with a distinct prompt pool. At test time, selecting the correct prompt pool for a given sample can be prone to errors. Such a misalignment may lead to the use of an inappropriate prompt pool, resulting in suboptimal or incorrect predictions. To mitigate this risk, we aim to improve the generalization capacity of each prompt pool, anticipating that it may encounter samples from tasks it was not originally designed for.
To achieve this, we propose augmenting the training objective for each prompt pool with representations derived from label descriptions. These representations capture the core attributes of each relation, thereby providing global information that helps reduce overfitting and improve model robustness. Previous studies have demonstrated that label descriptions can significantly benefit CRE tasks by providing additional context on relation types (Yang et al., 2020; Liu et al., 2022; Borchert et al., 2024; Luo et al., 2024). Notably, they offer a consistent, class-specific global perspective, enabling the model to focus on essential relational characteristics and avoid spurious features.
Specifically, we leverage the detailed definitions provided for each label, typically available in benchmarking datasets, which we refer to as raw descriptions. However, these raw descriptions may lack sufficient diversity and richness, which can introduce noise and instability that harm model performance. To address these limitations, we follow prior work (Thanh et al., 2025) and employ Gemini 1.5 (Team et al., 2023, 2024) to generate $D$ diverse, detailed, and illustrative descriptions for each relation. For each label, the raw description is fed into the large language model to guide the generation process. Our prompt template is shown in Table 1.
The generated descriptions are then encoded using a pre-trained BERT model to obtain their corresponding description representations. For task $t$, each relation $r\in\mathcal{R}_{t}$ is associated with a set of label description representations $Des_{r}=\{{\bm{d}}_{r,i}\in\mathbb{R}^{d}\ |\ i=1,\dots,D\}$. Given an input $\bm{x}$} with a ground-truth label $y$, our objective is to train prompts that align the input’s learned representation, $f_{\theta}({\bm{x}}_{p})$, with the description representations of the correct relation.
To achieve this, we pull $f_{\theta}({\bm{x}}_{p})$ closer to the description representations of the correct label while pushing it away from those of incorrect labels. We formalize this objective using a contrastive loss that enforces separation between the target and irrelevant labels:
@@ 4.2 Enhancing Training with Label Descriptions
where $Des_{y}$ is the set of description representations for the ground-truth label $y$. This loss function encourages high similarity between the input representation and its corresponding label descriptions while minimizing similarity with incorrect ones, thereby enhancing the model’s discriminative power. Consequently, incorporating label descriptions improves the stability and reliability of learned representations, leading to robust performance even when faced with previously unseen or misclassified samples.
@@
For each new task $\mathcal{T}_{t}$, we train a corresponding prompt pool $\mathcal{P}_{t}$. During each training step, we select $K$ prompts as previously described, and the resulting prompted embeddings, ${\bm{x}}_{p}$, are passed to the pre-trained transformer encoder $f_{\theta}$ and the final classifier $g_{\phi}$. The overall objective is to minimize the following loss function:
@@
where $\alpha$ and $\beta$ are hyperparameters. The first term the standard classification loss. Importantly, when training on task $t$, we only update the parameters of the current prompt pool $\mathcal{P}_{t}$ and the final classifier $g_{\phi}$. The pre-trained encoder and all previous prompt pools $\mathcal{P}_{1},\dots,\mathcal{P}_{t-1}$ remain frozen.
By combining task-specific prompt pools with label descriptions, our framework improves within-task prediction performance, a key component of the overall continual learning objective (see Equation (1)). Next, we introduce our strategy to enhance task-identity inference.
@@ 4.3 Cascade Voting for Task Prediction
Input: Voting results $\mathcal{V}^{0}({\bm{x}})$ and $\mathcal{V}^{1}({\bm{x}})$ of prompt pool $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, Maximum allowed experts $m$, Confidence scores $\{\text{Score}^{i}_{k}({\bm{x}})\ |\ 2\leq i\leq t-1,i\leq k\leq t\}$
@@
7:for $i\leftarrow start$ to endend do
8:  $\mathcal{V}^{i}({\bm{x}})=\underset{k}{\arg\min}\ \text{Score}^{i}_{k}({\bm{x}}),\ k\geq end$
@@
During training, the model is provided with a task-specific dataset and optimizes its corresponding prompt pool. However, during inference, a key challenge arises: the model no longer receives explicit task labels for incoming inputs. Consequently, it must autonomously determine which task-specific prompt pool to employ for generating accurate predictions.
To address this, HiDe-Prompt (Wang et al., 2024) introduces an additional MLP head $\hat{g}_{\psi}$, which serves as a task predictor. This head leverages the pre-trained unprompted feature $f_{\theta}({\bm{x}})$ to predict the task label. However, as discussed in Section 3, treating each task as a distinct class can be suboptimal because tasks are often defined by their order of appearance rather than by meaningful semantic differences. In response, Le et al. (2025a) increases the predictor’s output dimension to match the number of encountered relations, and then derives the task label from the classified relation. Despite this, this method still relies on pre-trained knowledge and overlooks the potential task-identification capabilities of the specialized experts $\mathcal{P}_{t}$.
In contrast, our approach leverages Cascade Voting (Zhou et al., 2024) to determine task identity. By discarding the MLP-based classifier head, Cascade Voting eliminates the need to train a task predictor. Instead, it employs a voting strategy in where the specialized experts in each prompt pool $\mathcal{P}_{t}$ is collaboratively determine the task identity for a given instance.
Formally, for an instance $\bm{x}$} of relation $r$, let ${\bm{z}}_{r}^{i}=f_{\theta,\mathcal{P}_{i}}({\bm{x}})\in\mathbb{R}^{d}$ be the representation obtained by passing $\bm{x}$} through the pre-trained transformer combined with prompt pool $\mathcal{P}_{i}$. We assume that during the $t$-th task, the class-conditional distribution of $\mathcal{D}^{i}_{r,t}=\{{\bm{z}}_{r}^{i}=f_{\theta,\mathcal{P}_{i}}({\bm{x}})|({\bm{x}},r)\in\mathcal{D}_{t}\}$ follows a multivariate Gaussian distribution $\mathcal{N}(\bm{\mu}^{i}_{r,t},\bm{\Sigma}^{i}_{t})$ for each prompt pool $\mathcal{P}_{i}$ and for all relations $r\in\mathcal{R}_{t}$. Based on this assumption, we use the maximum likelihood estimator to approximate the mean and covariance as follows:
@@
In continual relation extraction, the model is precluded from accessing data from previously encountered tasks. Consequently, it is impossible to establish a distribution for how prompt pool $\mathcal{P}_{i}$ affects an earlier task $j$ where $j<i$. To overcome it, we employ cascade voting during inference to determine the appropriate prompt pool for relation classification. This approach comprises two stages. First, each prompt pool, viewed as an expert specialized for a particular task, independently assesses and votes on the suitability for the given instance. Second, these votes are aggregated to identify the most suitable prompt pool for that instance.
Figure 4: Data flow diagram for the inference process. First, the task identity of the input $\bm{x}$} is inferred via cascade voting to determines the appropriate prompt pool. A query $q({\bm{x}})$, generated from the input is used to retrieve the prompt from this pool whose key most closely matches the query. The selected prompt is prepended to the input ${\bm{x}}_{e}$, creating the prompted input ${\bm{x}}_{p}$. This sequence is then fed into the BERT encoder. The output embeddings corresponding to the entities $E_{1}$ and $E_{2}$ are extracted and concatenated. Finally, the resulting concatenated embedding is passed to the relation classifier $g_{\phi}$ to predict the relation label $y$.
Voting Procedure for a Prompt Pool
The voting procedure for a prompt pool $\mathcal{P}_{i}$ begins by generating a feature representation ${\bm{z}}^{i}=f_{\theta,\mathcal{P}_{i}}({\bm{x}})$, by passing the input $\bm{x}$} through the pre-trained backbone $f_{\theta}$ combined with $\mathcal{P}_{i}$. Since each relation $r\in\mathcal{R}_{t}$ is associated with a pre-computed multivariate Gaussian distribution $\mathcal{N}(\bm{\mu}^{i}_{r,t},\bm{\Sigma}^{i}_{t})$, we calculate a voting score for each task $t\geqslant i$. In particular, the score for task $t$ is defined as the minimum Mahalanobis distance from ${\bm{z}}^{i}$ to the distribution $\{\mathcal{N}(\bm{\mu}^{i}_{r,t},\bm{\Sigma}^{i}_{t})\ |\ r\in\mathcal{R}_{t}\}$ of any relation within that task:
@@
Finally, the vote cast by $\mathcal{P}_{i}$ is the task identity that minimizes this score:
@@
Since each prompt pool’s valid prediction ranges can vary, we employ a cascade voting mechanism to mitigate these boundary mismatches. The pre-trained BERT model, $\mathcal{P}_{0}$, serves as the initial expert. We first consider the votes from $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$, as both are capable of predicting any subsequent task. If $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ concur on a task label, that label is immediately assigned to $\bm{x}$}.
If $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ yield conflicting predictions, we resolve the disagreement by aggregating votes from multiple experts. An expert $\mathcal{P}_{i}$ is permitted to predict task identities $j\geq i$. To maintain computational efficiency and a consistent number of voting experts, we define an index interval $[start,end]$ specifying the experts eligible to participate. For fairness, all experts within this interval make predictions over the task identity range $[end,t]$, where $t$ denotes the index of the current task. Since $\mathcal{P}_{0}$ can predict all task identities, we set $start=0$. To determine the upper bound endend while treating $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ symmetrically, we choose the smaller of their predicted task identities, ensuring that both predictions are included in the voting process. To prevent an excessive number of voting experts, we impose a threshold $m$ on the upper bound, and thus set $end=\min(\min(\mathcal{V}^{0},\mathcal{V}^{1}),m)$.
The task identity that receives the majority of votes is then assigned to $\bm{x}$}, and the corresponding prompt pool is selected for relation classification. This cascade voting strategy thus limits the number of prompt pools involved at each decision step and effectively handles inconsistencies in prediction ranges. A detailed algorithmic description is provided in Algorithm 1. As demonstrated in Section 5.3, our approach enhances task-identity prediction accuracy. This improvement boosts TII performance, which, according to Equation (1), leads to more effective continual relation extraction. For a comparison with the MLP-based approach, see C.
@@ 4.4 Mitigating Classification Bias with Generative Models
A key element of our framework is the MLP head relation classifier $g_{\phi}$. To mitigate catastrophic forgetting, we retain the representations of previously learned relations by deploying a generator that encodes their input distribution. Specifically, for each relation $r\in\mathcal{R}_{t}$, we use a Gaussian distribution $\mathcal{N}(\bm{\mu}^{t}_{r,t},\bm{\Sigma}^{t}_{t})$. The mean and covariance are computed from the prompt pool $\mathcal{P}_{t}$ of the current task, as described in Section 4.3.
@@
2:for $epoch\leftarrow 1$ to training_epochtraining\_epoch do
@@
7: Update $\hat{\mathcal{R}}_{t}\leftarrow\hat{\mathcal{R}}_{t-1}\cup\mathcal{R}_{t}$
@@
By using a prompt-based generative model, the representations of previously learned relations can be reconstructed and replayed for subsequent tasks, thereby preserving essential relational knowledge. Storage requirements for each relation are minimal, as only a mean vector and a covariance matrix are stored. Furthermore, this design effectively addresses catastrophic forgetting by eliminating the need to store observed instances explicitly. Instead, the generative models capture and reproduce the distributional characteristics of each relation’s representations, enabling the model to simulate previously encountered relations without directly accessing historical data. This approach not only enhances computational efficiency but also ensures compliance with privacy constraints, as raw data is never retained.
These generative models produce synthetic representations ${\bm{z}}_{r}$ for each relation $r$, which are then used to train the relation classifier. The MLP classifier $g_{\phi}$ is trained by minimizing the following cross-entropy loss:
$\mathcal{L}(\phi)=\sum_{i=1}^{t}\sum_{r\in\mathcal{R}_{i}}\sum_{{\bm{z}}_{r}\sim\mathcal{N}(\bm{\mu}^{t}_{r,t},\bm{\Sigma}^{t}_{t})}-\mathrm{log}\frac{\exp(g_{\phi}({\bm{z}}_{r})[r])}{\sum_{r^{\prime}\in\widehat{\mathcal{R}}_{t}}\exp(g_{\phi}({\bm{z}}_{r})[r^{\prime}])}.$ (17)
The complete training procedure is detailed in Algorithm 2, and the data flow for the inference process is illustrated in Figure 4.
@@
5 Experiments
5.1 Experimental Settings
Datasets
To ensure consistency and fairness in our evaluation, we follow the protocols established by Zhou et al. (2024); Le et al. (2025a) and assess the performance of WAVE++ on two widely used CRE datasets:
  • 1.
    TACRED (Zhang et al., 2017) comprises 106,264 instances across 41 relation types. Following the experimental setup of Cui et al. (2021), we partition this dataset into 10 distinct subsets. The number of training samples of each relation is limited to 320, while the number of test samples of each relation is limited to 40.
  • 2.
    FewRel (Han et al., 2018) contains 56,000 instances across 80 relation types. Following the setup of Wang et al. (2019), we divide this dataset into 10 distinct subsets.
Baseline Models
We compare WAVE++ with several prompt-based methods, including L2P (Wang et al., 2022b), HiDe-Prompt (Wang et al., 2024), EoE (Zhou et al., 2024) and WAVE-CRE (Le et al., 2025a). Notably, WAVE-CRE (Le et al., 2025a) is a simplified version of our proposed WAVE++. While WAVE-CRE also employs task-specific prompt pools, WAVE++ extends this strategy by integrating label descriptions to better capture relational information, and replacing the MLP classifier with cascade voting for task identity prediction. We further contrast our approach with traditional rehearsal-based methods: RP-CRE (Cui et al., 2021), ACA (Wang et al., 2022b), CRL (Zhao et al., 2022), CDec (Xia et al., 2023), CEAR (Zhao et al., 2023), RationaleCL (Xiong et al., 2023), CREST (Le et al., 2024b), and DP-CRE (Huang et al., 2024). For all methods, we utilize a pre-trained BERT encoder (Devlin, 2018) as the backbone architecture. We report model performance in terms of average accuracy across five distinct random runs.